x+2y=6
y=2x+3
use the substitution method

Answer by stanbon(57250)

Since y=2x+3 substitute into the 1st equation to get:
x+2(2x+3)=6
x+4x+6=6
5x=0
x=0
Substitute that backinto either of the original equations to get:
y=2(0)+3
y=3
Solution is x=0;y=3
Check that in the other original equation.
Cheers,
Stan H.